# Multi-person Physics-based Pose Estimation for Combat Sports

### v2

@@ 2 Related Work
Recent physics-based approaches to single- and multi-person 3D pose estimation integrate differentiable physics, biomechanics, and reinforcement learning to mitigate common artifacts such as foot sliding, body inter-penetration, and collisions [23, 61, 20, 59, 17, 16, 46, 48, 14, 38]. Differentiable physics-based methods explicitly enforce biomechanical constraints by jointly optimizing kinematic and dynamic parameters, ensuring physically realistic motion even under challenging athletic activities [17, 16, 38]. Similarly, reinforcement learning frameworks enable simulated agents to refine pose predictions through physics-informed rewards, reducing artifacts like floating body parts and penetration [61, 20]. Incorporation of physics modeling has further improved motion plausibility even when using sparse IMU signals as input, particularly by preventing foot sliding and stabilizing the motion [1, 59, 46]. Recent advances also address multi-person scenarios by enforcing mutual constraints, ensuring spatial coherence and collision-free motion [48, 14]. These methods often benefit from human-scene interaction cues, leveraging ground support and physical laws to promote robust pose estimation in dense crowds. Overall, the integration of physics-based modeling enhances realism and robustness across diverse applications, establishing a new standard for performance in both single- and multi-person pose reconstruction tasks. The incorporation of biomechanical constraints—as shown by Saleem et al.’s BioPose [39]—ensures realistic and immersive human motion. Moreover, reinforcement learning approaches, such as the imitation-based method proposed by Peng et al. [36], contribute to generating physically coherent human animations; collectively, these diverse methodologies offer robust and realistic multi-person motion estimation suitable for sports analysis, crowd dynamics [37, 40, 27, 9, 29].
@@
Multi-person 3D pose estimation in sports poses significant challenges due to fast and unpredictable movements, frequent occlusions, and the need for fine-grained motion capture under high-intensity conditions [5]; early methods relied on controlled environments with multi-camera setups or marker-based motion capture systems that, while accurate, proved costly and inflexible for real-world scenarios. Recent advances have shifted toward leveraging monocular or multi-view video feeds in unconstrained settings to enable applications such as player tracking, action recognition, and performance analysis in sports like basketball, soccer [24], and tennis [49]. To address issues arising from frequent occlusions and complex player interactions, many approaches now integrate spatio-temporal modules or graph-based methods to capture both short-term and long-term dynamics in high-speed sports sequences [35, 56], while Transformer-based architectures have shown promise in effectively modeling global context in dense player formations [43]. The emergence of large-scale sports datasets with rich annotations further facilitates improved generalization under domain-specific constraints [22], spurring the development of specialized datasets and systems such as SportsPose [22], which features markerless motion capture of five high-speed sports validated against commercial marker-based systems, and AthletePose3D [58], which captures extreme dynamics across 12 sport motion types and approximately 1.3 million frames to achieve a reduction in pose error by nearly 69% (from 214 mm to 65 mm). For team sports, WorldPose [24] provides a large-scale multi-person dataset from the 2022 FIFA World Cup that captures 3D poses and trajectories of entire soccer teams using a static multi-camera setup, while tailored applications like Reconstructing NBA Players [64] demonstrate the integration of computer vision with domain-specific data to achieve high-resolution 3D mesh reconstruction from single broadcast images despite challenges such as complex poses, motion blur, and occlusions. In dense crowd scenarios, robust multi-view geometry frameworks that combine probabilistic reasoning and cross-view matching have proven effective [6]. In this work, we focus on using multi-person physics-based pose estimation of elite boxers captured during sparring sessions to robustly recover detailed 3D kinematics under challenging occlusion and rapid motion conditions. Our approach uniquely integrates physics simulation with advanced deep learning, enabling precise biomechanical analysis and offering new insights for performance optimization in competitive boxing.
@@ 3 Methodology
Figure LABEL:fig:3d illustrates the overall architecture of our proposed framework, comprising four key stages: multi-view tracking, weighted triangulation, kinematic optimization, and dynamics-based trajectory refinement. Each stage is detailed in Sections 3.1 through 3.4.
@@ 3.1 Multi-frame Multi-view Tracking IDs
Figure 3: Illustration of the epipolar geometry distance for a pair of cross view 2D positions
@@ 3.4 Multi-person Dynamic Optimization
Figure 4: Illustration of SMPL to physics-based humanoid conversion for varying body shapes

### v3

@@ 2 Related Work
Recent physics-based approaches to single- and multi-person 3D pose estimation integrate differentiable physics, biomechanics, and reinforcement learning to mitigate common artifacts such as foot sliding, body inter-penetration, and collisions [25, 61, 22, 58, 20, 19, 48, 50, 16, 60]. Differentiable physics-based methods explicitly enforce biomechanical constraints by jointly optimizing kinematic and dynamic parameters, ensuring physically realistic motion even under challenging athletic activities [20, 19, 60]. Similarly, reinforcement learning frameworks enable simulated agents to refine pose predictions through physics-informed rewards, reducing artifacts like floating body parts and penetration [61, 22]. Incorporation of physics modeling has further improved motion plausibility even when using sparse IMU signals as input, particularly by preventing foot sliding and stabilizing the motion [2, 58, 48]. Recent advances also address multi-person scenarios by enforcing mutual constraints, ensuring spatial coherence and collision-free motion [50, 16]. These methods often benefit from human-scene interaction cues, leveraging ground support and physical laws to promote robust pose estimation in dense crowds. Overall, the integration of physics-based modeling enhances realism and robustness across diverse applications, establishing a new standard for performance in both single- and multi-person pose reconstruction tasks. The incorporation of biomechanical constraints—as shown by Saleem et al.’s BioPose [40]—ensures realistic and immersive human motion. Moreover, reinforcement learning approaches, such as the imitation-based method proposed by Peng et al. [38], contribute to generating physically coherent human animations; collectively, these diverse methodologies offer robust and realistic multi-person motion estimation suitable for sports analysis, crowd dynamics [60, 42, 29].
@@
Multi-person 3D pose estimation in sports poses significant challenges due to fast and unpredictable movements, frequent occlusions, and the need for fine-grained motion capture under high-intensity conditions [6]; early methods relied on controlled environments with multi-camera setups or marker-based motion capture systems that, while accurate, proved costly and inflexible for real-world scenarios. Recent advances have shifted toward leveraging monocular or multi-view video feeds in unconstrained settings to enable applications such as player tracking, action recognition, and performance analysis in sports like basketball, soccer [7, 26], and tennis [1]. To address issues arising from frequent occlusions and complex player interactions, many approaches now integrate spatio-temporal modules or graph-based methods to capture both short-term and long-term dynamics in high-speed sports sequences [39, 45], while Transformer-based architectures have shown promise in effectively modeling global context in dense player formations [41]. The emergence of large-scale sports datasets with rich annotations further facilitates improved generalization under domain-specific constraints [30], spurring the development of specialized datasets and systems such as SportsPose [24], which features markerless motion capture of five high-speed sports validated against commercial marker-based systems, and AthletePose3D [57], which captures extreme dynamics across 12 sport motion types and approximately 1.3 million frames to achieve a reduction in pose error by nearly 69% (from 214 mm to 65 mm). For team sports, WorldPose [26] provides a large-scale multi-person dataset from the 2022 FIFA World Cup that captures 3D poses and trajectories of entire soccer teams using a static multi-camera setup, while tailored applications like Reconstructing NBA Players [64] demonstrate the integration of computer vision with domain-specific data to achieve high-resolution 3D mesh reconstruction from single broadcast images despite challenges such as complex poses, motion blur, and occlusions. In dense crowd scenarios, robust multi-view geometry frameworks that combine probabilistic reasoning and cross-view matching have proven effective [8]. In this work, we focus on using multi-person physics-based pose estimation of elite boxers captured during sparring sessions to robustly recover detailed 3D kinematics under challenging occlusion and rapid motion conditions. Our approach uniquely integrates physics simulation with advanced deep learning, enabling precise biomechanical analysis and offering new insights for performance optimization in competitive boxing.
@@ 3 Methodology
Figure 2 illustrates the overall architecture of our proposed framework, comprising four key stages: multi-view tracking, weighted triangulation, kinematic optimization, and dynamics-based trajectory refinement. Each stage is detailed in Sections 3.1 through 3.4.
[…]
@@ 3.1 Multi-frame Multi-view Tracking IDs
[…]
Figure 3: Illustration of the epipolar geometry distance for a pair of cross view 2D positions.
@@ 3.4 Multi-person Dynamic Optimization
Figure 4: Illustration of SMPL to physics-based humanoid conversion for varying body shapes.